# Kernel Ridge Regression for conformer ensembles made easy with Structured Orthogonal Random Features

### v1

@@ Abstract
A computationally efficient protocol for machine learning in chemical space using Boltzmann ensembles of conformers as input is proposed; the method is based on rewriting Kernel Ridge Regression expressions in terms of Structured Orthogonal Random Features, yielding physics-motivated trigonometric neural networks. The resulting method is tested on experimental datasets of oxidation potentials in acetonitrile and hydration energies, using several popular molecular representations to demonstrate the method’s flexibility. Despite only using computationally cheap forcefield calculations for conformer generation, we observe systematic decrease of machine learning error with increased training set size in all cases, with experimental accuracy reached after training on hundreds of molecules and prediction errors being comparable to state-of-the-art machine learning approaches. We also present novel versions of Huber and LogCosh loss functions that made hyperparameter optimization of the new approach more convenient.
@@ 2.1 Kernel Ridge Regression in space of molecular conformers
We will now outline three strategies for representing molecular conformers and defining the kernel function $k$ in terms of those representations which were used in this work. The first strategy is to represent each molecule $\mathrm{A}$ with a vector $\vec{X}^{(\mathrm{A})}$ and then calculate the kernel function $k(\mathrm{A},\mathrm{B})$ using the (global) Gaussian kernel function
@@
where ||$\ldots$|| … | is a vector’s $l^{2}$ norm and $\sigma$ is a hyperparameter. Examples of vector molecular representations considered in this work are Coulomb Matrix (CM) [37, 38] and Spectrum of London and Axilrod-Teller-Muto potential (SLATM) [4].
@@
where $N_{\mathrm{a}}^{(\mathrm{A})}$ and $N_{\mathrm{a}}^{(\mathrm{B})}$ are number of atoms in molecules $\mathrm{A}$ and $\mathrm{B}$, $n_{j^{\prime}}^{(A)}$ and $n_{j^{\prime\prime}}^{(B)}$ are nuclear charges of atoms $j^{\prime}$ and $j^{\prime\prime}$ of molecules $\mathrm{A}$ and $\mathrm{B}$, while $\vec{X}_{j^{\prime}}^{(\mathrm{A})}$ and $\vec{X}_{j^{\prime\prime}}^{(\mathrm{B})}$ are the corresponding atomic representation vectors. $k_{\mathrm{local}}$ is particularly useful for ML of extensive molecular quantities (e.g. total potential energy); an example of a representation that can be used with $k_{\mathrm{local}}$ is Smooth Overlap of Atomic Orbitals (SOAP) descriptors [39, 3]. Another popular kernel expression is
@@
where $\delta$ is the Kroneker delta function. Examples of representations used with $k_{\mathrm{dlocal}}$ considered in this work are revised version of Faber-Christensen-Huang-Lilienfeld representation (FCHL19) [40, 41], atomic version of SLATM (aSLATM) [4], and convolutional version of Many Body Distribution Functional (cMBDF) [42, 43].
The third kernel expression we will consider is specific to the FJK representation [44] (named so due to Fock $\mathbf{F}$, Coulomb $\mathbf{J}$, and exchange $\mathbf{K}$ used to generate it). The representation of molecule $\mathrm{A}$ takes results of cheap Hartree Fock calculations, among them properties of localized orbitals $\psi_{o}^{(\mathrm{A})}$ ($o=1,\ldots,N_{\mathrm{orb}}^{(\mathrm{A})}$, where $N_{\mathrm{orb}}^{(\mathrm{A})}$ is the number of occupied orbitals in molecule $\mathrm{A}$), then assigns a weight $w_{j}(\psi_{o}^{(\mathrm{A})})$ and vector representation $\vec{v}_{j}(\psi_{o}^{(\mathrm{A})})$ to each contribution to the localized orbital from atomic orbitals centered on an atom with index $j$ ($j\in 1,\ldots,N_{\mathrm{a}}^{(\mathrm{A})}$). The corresponding expression for kernel function between molecules $\mathrm{A}$ and $\mathrm{B}$ can be written as
@@
where $k_{\mathrm{FJK}}^{\mathrm{lin}}$ and $k_{\mathrm{FJK}}^{\mathrm{Gauss}}$ are linear and Gaussian kernel functions for localized orbitals, $\sigma_{\mathrm{global}}$ is a hyperparameter, $\vec{\sigma}$ is a vector of hyperparameters of the same dimensionality as $\vec{v}_{j}(\psi_{o}^{(\mathrm{A})})$ (the implied division by this vector is element-wise), and $\odot$ stands for element-wise multiplication. [^1] The definition of $k_{\mathrm{FJK}}^{\mathrm{lin}}$ (8) differs from the original one in Ref [44] by rescaling $\vec{\sigma}$ by a factor of $\sqrt{2}$. Note that the number of weights $w_{j^{\prime}}(\psi_{o^{\prime}}^{(\mathrm{A})})$ and $w_{j^{\prime\prime}}(\psi_{o^{\prime\prime}}^{(\mathrm{B})})$ that are not exactly zero does not grow with system size, meaning the same can be said about the computational cost of evaluating $k_{\mathrm{FJK}}^{\mathrm{lin}}$ (8). $\vec{v}(\psi_{o}^{(\mathrm{A})})$ consists of entries corresponding to momentum distribution, $\mathbf{F}$, $\mathbf{J}$, and $\mathbf{K}$ matrices; in this work these four different types of entries were given a separate $\sigma$ value each (which differs from the $\vec{\sigma}$ used in Ref. [44]).
@@ 2.2 Rewriting kernels in terms of Structured Orthogonal Random Features
$\vec{\mathcal{F}}(\vec{X}^{(\mathrm{A})},\mathrm{D},\upphi,\sigma)\cdot\vec{\mathcal{F}}(\vec{X}^{(\mathrm{B})},\mathrm{D},\upphi,\sigma)\approx k_{\mathrm{global}}(\mathrm{A},\mathrm{B}).$ (14)
We note at this point that all the applications discussed in this work in the context of SORF can be analogously performed with other procedures [19, 20] for generating feature vectors that approximate $k_{\mathrm{global}}$ analogously to Equation (14). However, in this work we used SORF due to its balance of good numerical performance and algorithmic simplicity.
@@
$\mathcal{S}\mathcal{F}(\mathbf{X}^{(\mathrm{A})},\mathrm{D},\upphi,\sigma):=\sum_{j=1}^{N_{\mathrm{a}}}\mathcal{F}(\vec{X}_{j}^{(\mathrm{A})},\mathrm{D},\upphi,\sigma).$ (15)
@@
$\mathcal{S}\mathcal{T}\mathcal{F}(\mathbf{X}^{(\mathrm{A})},\sigma):=\sum_{j}^{N_{\mathrm{a}}}\mathcal{F}[\vec{X}_{j}^{(\mathrm{A})},\mathrm{D}(n_{j}^{(\mathrm{A})}),\upphi(n_{j}^{(\mathrm{A})}),\sigma],$ (16)
@@
$k_{\mathrm{FJK}}^{\mathrm{pair}}(\mathrm{A}_{1\rightarrow 2},\mathrm{B}_{1\rightarrow 2})\approx[\mathcal{S}\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}(\mathrm{A}_{2})-\mathcal{S}\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}(\mathrm{A}_{1})][\mathcal{S}\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}(\mathrm{B}_{2})-\mathcal{S}\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}(\mathrm{B}_{1})].$ (20)
@@ 2.3 Applications to ensembles of conformers
For intensive properties, we aim to create models whose prediction is invariant w.r.t. “molecule multiplication”, i.e. creating several copies of a molecule far away from each other. For global representations the requirement again cannot be satisfied implicitly, thus simple $\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}$ was used. For local representations $\mathcal{N}\mathcal{S}\mathcal{F}$ or $\mathcal{N}\mathcal{S}\mathcal{T}\mathcal{F}$ generates a feature vector for a conformer invariant w.r.t. molecule multiplication; treating it as a global representation vector yields $\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}\mathcal{N}\mathcal{S}\mathcal{F}$ and $\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}\mathcal{N}\mathcal{S}\mathcal{T}\mathcal{F}$ as the final layer combinations depending on whether $k_{\mathrm{local}}$ or $k_{\mathrm{dlocal}}$ is reproduced.
@@
where $\rho_{\mathrm{cut}}$ is a parameter chosen by the user; note that $\vec{w}^{\mathrm{cb}}$ changes continuously with $w^{\mathrm{b}}$. [^2] This is the same procedure as the one used in FJK to avoid processing irrelevant contributions to localized orbitals; the only difference is the normalization factor. The procedure was introduced to avoid calculating features of conformers whose contribution is irrelevant. Secondly, our algorithm’s implementation performs conformer generation not once, but several times to potentially account for statistical noise. Though we did not use the latter option and conformers were generated only once for each molecule, code-wise we were still performing “summation over a single conformer generation.” As a result in the actual scripts used in this work each $\mathcal{W}$ layer corresponding to summation over conformers (see Table 1) was preceded by an additional $\mathcal{S}$ layer corresponding to summation over conformer generations.
@@ 3 Numerical details
For each learning curve presented in this work we divided a dataset into a training set (approximately 80% of the total number of points, making it 474 for LES and 514 for FreeSolv) and a test set. For each training size we picked a random subset of the training set, used it to optimize hyperparameters as described in A, trained the model using these hyperparameters, and then calculated Mean Absolute Error (MAE) of model’s predictions for the test set molecules. For each dataset and method the procedure was repeated 4 times, providing both the average MAEs and their dispersions presented in Section 4.
@@
By default we ran calculations with $N_{\mathrm{f}}=32768$ at all $\mathcal{F}$ layers, since the value had performed well in Ref. [21], and chose $N_{\mathrm{init}}$ as the smallest power of two integer greater or equal to representation vector size, padding the “extra” values of the vector with zeros. This could not be done for SLATM, aSLATM, and (in the case of LES) SOAP since their representation vectors corresponded to $N_{\mathrm{init}}$ larger than $32768$. For SLATM, the issue was resolved by projecting all representation vectors onto linear space spawned by representation vectors of conformers in the training set, and then appending to the resulting vectors the $l^{2}$ distance between the projected and non-projected representation vector; note that using thus defined new representation vectors would not affect training and predictions of a KRR model. The procedure decreased $N_{\mathrm{init}}$ to $4096$ for SLATM, but could not be used for aSLATM and SOAP due to large number of atomic environments in the training set. Ref. [21] used Principal Component Analysis to fit atomic representations into a small enough $N_{\mathrm{init}}$, but we wanted to avoid that in order to accurately observe prediction errors of aSLATM and SOAP relative to other methods. Therefore, for aSLATM we used $N_{\mathrm{init}}=131072$, for SOAP we used $N_{\mathrm{init}}=65536$ for LES and $N_{\mathrm{init}}=32768$ for FreeSolv.
@@
We generated molecular conformers with the algorithm from Ref. [46], as implemented in the Morfeus package [47], at $T=298.15\,\mathrm{K}$ with the Merck molecular force field (MMFF94) [48, 49, 50, 51, 52, 53, 54] as implemented [55] in RDKit [56]; unimportant conformers were cut off with $\rho_{\mathrm{cut}}=0.05$. Morfeus calculations require a parameter $N_{\mathrm{conf}}^{\mathrm{Morfeus}}$ determining the number of conformers the package attempts to generate; we checked convergence of the $\mathbf{Z}\mathbf{Z}^{\mathbf{T}}$ matrix generated for one of the 4 training sets chosen for the learning curves with cMBDF representations by scanning $N_{\mathrm{conf}}^{\mathrm{Morfeus}}=8,\ldots,1024$ and observed that the values were converged at $N_{\mathrm{conf}}^{\mathrm{Morfeus}}=32$ for both LES and FreeSolv, hence $N_{\mathrm{conf}}^{\mathrm{Morfeus}}=32$ was used throughout the numerical results. In both cases, the $\sigma$ and $\vec{\sigma}$ hyperparameters used in $\mathcal{F}$ layers were taken to be the initial guess defined in A. We used the same convergence criterion for optimizing $N_{\mathrm{tr}}$ by calculating a training set’s $\mathbf{Z}\mathbf{Z}^{\mathbf{T}}$ $N_{\mathrm{tr}}=1,2,\ldots,6,7$ for each representation and combination of MSORF layers considered. We settled on using $N_{\mathrm{tr}}=3$ throughout the work (which is also consistent with observations in Refs. [17] and [21]).
@@
Almost all representation functions used in this work were calculated using the QML2 code [57], the exception being SOAP which was calculated with DScribe package [58, 59]. While aSLATM, SLATM, FCHL19, and cMBDF were generated using default parameters present in QML2, SOAP was generated using $r_{\mathrm{cut}}^{\mathrm{SOAP}}=9.53$, $n_{\mathrm{max}}^{\mathrm{SOAP}}=8$, and $l_{\mathrm{max}}^{\mathrm{SOAP}}=8$, which are the parameters used in Ref. [60] for ML on FreeSolv with SOAP descriptors and KRR. FJK representations were generated from ab initio calculations performed with the PySCF package [61, 62, 63]. While the original paper where FJK was proposed [44] used STO-3G basis set [64, 65] and intrinsic bond orbitals [66], we found both to be ill-suited for the datasets considered in this work: Hartree-Fock STO-3G calculations sometimes failed to converge for charged molecules (which we wanted to consider as part of Slater determinant pairs), while using intrinsic bond orbitals in PySCF is complicated with some elements. Taking inspiration from previous work on using Huckel calculations for ML [7, 67], we switched to Huckel calculations with 6-311G basis [68, 69, 70, 71] and Boys localized orbitals [72] to generate FJK representations. Presence of solvent was simulated with Solvation Model based on Density (SMD) [73], an ab initio approach that has proven to be accurate in predicting experimental hydration energies [12]. When referring to “single Slater determinant” FJK we will mean representations calculated from ground state Huckel calculations for conformers in solvent (acetonitrile for LES and water for FreeSolv); for “pair Slater determinant” FJK for LES we took ground states of the molecule at charges 0 and 1 (both solvated in acetonitrile), while for FreeSolv we took ground states of the molecule in vacuum and solvated in water.
@@
While the full procedure used to optimize hyperparameters is described in A, here we note that apart from optimizing $\lambda$ value and different $\sigma$ hyperparameters, we also considered optimizing quantity of interest by shifting it by a constant if it is intensive or by a stochiometry-dependent shift [see Eq. (29)] if it is extensive. Such optimization would mean, for example, that ML of oxidation potentials is invariant w.r.t. the choice of the reference electrode, while ML of a compound’s free energy of formation and atomization are equivalent since they only differ in reference state for different elements. However, the procedure did not change the results significantly for LES, while for FreeSolv it became numerically unstable for smaller dataset sizes (probably due to the large number of atom types present, leading to a large number of addtional shift-related hyperparameters). Therefore for LES in Subsection 4.1 we present results obtained while applying the quantity shifts, while for FreeSolv in Subsection 4.2 the results were obtained without them. Learning curves without quantity shifts for LES and with them for FreeSolv were left for Supplementary Data.

### v2

@@ Abstract
A computationally efficient protocol for machine learning in chemical space using Boltzmann ensembles of conformers as input is proposed; the method is based on rewriting Kernel Ridge Regression expressions in terms of Structured Orthogonal Random Features, yielding physics-motivated trigonometric neural networks. To evaluate the method’s utility for materials discovery, we test it on experimental datasets of two quantities related to battery electrolyte design, namely oxidation potentials in acetonitrile and hydration energies, using several popular molecular representations to demonstrate the method’s flexibility. Despite only using computationally cheap forcefield calculations for conformer generation, we observe systematic decrease of machine learning error with increased training set size in all cases, with experimental accuracy reached after training on hundreds of molecules and prediction errors being comparable to state-of-the-art machine learning approaches. We also present novel versions of Huber and LogCosh loss functions that made hyperparameter optimization of the new approach more convenient.
@@ 2.1 Kernel Ridge Regression in space of molecular conformers
We will now outline three strategies for representing molecular conformers and defining the kernel function $k$ in terms of those representations that were used in this work. The first strategy is to represent each molecule $\mathrm{A}$ with a vector $\vec{X}^{(\mathrm{A})}$ and then calculate the kernel function $k(\mathrm{A},\mathrm{B})$ using the (global) Gaussian kernel function
@@
where ||$\ldots$|| … | is a vector’s $l^{2}$ norm and $\sigma$ is a hyperparameter. Examples of such molecular representations considered in this work are Coulomb Matrix (CM) [37, 38] and Spectrum of London and Axilrod-Teller-Muto potential (SLATM) [4].
@@
where $\vec{X}_{j^{\prime}}^{(\mathrm{A})}$ and $\vec{X}_{j^{\prime\prime}}^{(\mathrm{B})}$ are the atomic representation vectors of atoms with indices $j^{\prime}$ and $j^{\prime\prime}$ in molecules $\mathrm{A}$ and $\mathrm{B}$, while $N_{\mathrm{a}}^{(\mathrm{A})}$ and $N_{\mathrm{a}}^{(\mathrm{B})}$ are numbers of atoms in molecules $\mathrm{A}$ and $\mathrm{B}$. $k_{\mathrm{local}}$ is particularly useful for ML of extensive molecular quantities (e.g. total potential energy); an example of a representation that can be used with $k_{\mathrm{local}}$ is Smooth Overlap of Atomic Orbitals (SOAP) descriptors [39, 3]. The kernel expression can be modified as
@@
where $n_{j^{\prime}}^{(A)}$ and $n_{j^{\prime\prime}}^{(B)}$ are nuclear charges of atoms $j^{\prime}$ and $j^{\prime\prime}$ in molecules $\mathrm{A}$ and $\mathrm{B}$, and $\delta$ is the Kroneker delta function. Examples of representations used with $k_{\mathrm{dlocal}}$ considered in this work are revised version of Faber-Christensen-Huang-Lilienfeld representation (FCHL19) [40, 41], atomic version of SLATM (aSLATM) [4], and convolutional version of Many Body Distribution Functional (cMBDF) [42, 43].
The third strategy we will consider is specific to the FJK representation [44] (named so due to Fock $\mathbf{F}$, Coulomb $\mathbf{J}$, and exchange $\mathbf{K}$ used to generate it). The representation of molecule $\mathrm{A}$ takes results of cheap Hartree Fock calculations, among them properties of localized orbitals $\psi_{o}^{(\mathrm{A})}$ ($o=1,\ldots,N_{\mathrm{orb}}^{(\mathrm{A})}$, where $N_{\mathrm{orb}}^{(\mathrm{A})}$ is the number of occupied orbitals in molecule $\mathrm{A}$), then assigns a weight $w_{j}(\psi_{o}^{(\mathrm{A})})$ and vector representation $\vec{v}_{j}(\psi_{o}^{(\mathrm{A})})$ to each contribution to the localized orbital from atomic orbitals centered on an atom with index $j$ ($j\in 1,\ldots,N_{\mathrm{a}}^{(\mathrm{A})}$). The corresponding expression for kernel function between molecules $\mathrm{A}$ and $\mathrm{B}$ can be written as
@@
where $k_{\mathrm{FJK}}^{\mathrm{lin}}$ and $k_{\mathrm{FJK}}^{\mathrm{Gauss}}$ are linear and Gaussian kernel functions for localized orbitals, $\sigma_{\mathrm{global}}$ is a hyperparameter, $\vec{\sigma}$ is a vector of hyperparameters of the same dimensionality as $\vec{v}_{j}(\psi_{o}^{(\mathrm{A})})$ (the implied division by this vector is element-wise), and $\odot$ stands for element-wise multiplication. [^1] The definition of $k_{\mathrm{FJK}}^{\mathrm{lin}}$ (8) differs from the original one in Ref [44] by rescaling $\vec{\sigma}$ by a factor of $\sqrt{2}$. The number of weights $w_{j^{\prime}}(\psi_{o^{\prime}}^{(\mathrm{A})})$ and $w_{j^{\prime\prime}}(\psi_{o^{\prime\prime}}^{(\mathrm{B})})$ that are not exactly zero does not grow with system size, meaning the same can be said about the computational cost of evaluating $k_{\mathrm{FJK}}^{\mathrm{lin}}$ (8). $\vec{v}(\psi_{o}^{(\mathrm{A})})$ consists of entries corresponding to momentum distribution, $\mathbf{F}$, $\mathbf{J}$, and $\mathbf{K}$ matrices; in this work these four different types of entries were given a separate $\sigma$ value each (which differs from the $\vec{\sigma}$ used in Ref. [44]).
@@ 2.2 Rewriting kernels in terms of Structured Orthogonal Random Features
$\vec{\mathcal{F}}(\vec{X}^{(\mathrm{A})},\mathrm{D},\upphi,\sigma)\cdot\vec{\mathcal{F}}(\vec{X}^{(\mathrm{B})},\mathrm{D},\upphi,\sigma)\approx k_{\mathrm{global}}(\mathrm{A},\mathrm{B}),$ (14)
with the approximation introducing a statistical error which can be decreased by increasing $N_{\mathrm{stack}}$ and a systematic bias which can be decreased by increasing $N_{\mathrm{init}}$ [17] (e.g. by padding the input vector with zeros). We note at this point that all the applications discussed in this work in the context of SORF can be analogously performed with other procedures [19, 20] for generating feature vectors that approximate $k_{\mathrm{global}}$ analogously to Equation (14). However, in this work we used SORF due to its balance of good numerical performance and algorithmic simplicity.
@@
$\mathcal{S}\mathcal{F}(\mathbf{X}^{(\mathrm{A})},\mathrm{D},\upphi,\sigma):=\sum_{j=1}^{N_{\mathrm{a}}^{(\mathrm{A})}}\mathcal{F}(\vec{X}_{j}^{(\mathrm{A})},\mathrm{D},\upphi,\sigma).$ (15)
@@
$\mathcal{S}\mathcal{T}\mathcal{F}(\mathbf{X}^{(\mathrm{A})},\sigma):=\sum_{j}^{N_{\mathrm{a}}^{(\mathrm{A})}}\mathcal{F}[\vec{X}_{j}^{(\mathrm{A})},\mathrm{D}(n_{j}^{(\mathrm{A})}),\upphi(n_{j}^{(\mathrm{A})}),\sigma],$ (16)
@@
$k_{\mathrm{FJK}}^{\mathrm{pair}}(\mathrm{A}_{1\rightarrow 2},\mathrm{B}_{1\rightarrow 2})\approx[\mathcal{S}\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}(\mathrm{A}_{2})-\mathcal{S}\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}(\mathrm{A}_{1})]\cdot[\mathcal{S}\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}(\mathrm{B}_{2})-\mathcal{S}\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}(\mathrm{B}_{1})].$ (20)
@@ 2.3 Applications to ensembles of conformers
For intensive properties, we aim to create models whose prediction is invariant w.r.t. “molecule multiplication”, i.e. creating several copies of a molecule far away from each other. For global representations the requirement again cannot be satisfied implicitly, thus simple $\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}$ was used. For local representations $\mathcal{N}\mathcal{S}\mathcal{F}$ or $\mathcal{N}\mathcal{S}\mathcal{T}\mathcal{F}$ generates a feature vector for a conformer invariant w.r.t. molecule multiplication; treating it as a global representation vector yields $\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}\mathcal{N}\mathcal{S}\mathcal{F}$ and $\mathcal{F}\mathcal{N}\mathcal{W}\mathcal{F}\mathcal{N}\mathcal{S}\mathcal{T}\mathcal{F}$ as the final layer combinations.
@@
where $\rho_{\mathrm{cut}}$ is a parameter chosen by the user; note that $\vec{w}^{\mathrm{cb}}$ changes continuously with $w^{\mathrm{b}}$. [^2] This is the same procedure as the one used in FJK to avoid processing irrelevant contributions to localized orbitals; the only difference is the normalization factor. The procedure was introduced to avoid calculating features of conformers whose contribution is irrelevant. Secondly, our algorithm’s implementation allows performing conformer generation not once, but several times to potentially account for statistical noise. Though we did not use the latter option and conformers were generated only once for each molecule, code-wise we were still performing “summation over a single conformer generation.” As a result in the actual scripts used in this work each $\mathcal{W}$ layer corresponding to summation over conformers (see Table 1) was preceded by an additional $\mathcal{S}$ layer corresponding to summation over conformer generations.
@@ 3 Numerical details
For each learning curve presented in this work we divided a dataset into a training set (approximately 80% of the total number of points, making it 474 for LES and 514 for FreeSolv) and a test set. For each training size we picked a random subset of the training set, used it to optimize hyperparameters as described in A, trained the model using these hyperparameters, and then calculated Mean Absolute Error (MAE) of model’s predictions for the test set molecules. For each dataset and method the procedure was repeated 4 times, providing both the average MAEs and their dispersions presented in Section 4. Since the datasets used in this work a collections of literature data gathered over inhomogeneous experimental reports, we consider it more natural to measure ML model accuracy in terms of MAE rather than Root Mean Square Error (RMSE), which is more punishing towards outliers. However, for future reference we will still present RMSEs observed for largest training set sizes.
@@
By default we ran calculations with $N_{\mathrm{f}}=32768$ at all $\mathcal{F}$ layers, since the value had performed well in Ref. [21], and chose $N_{\mathrm{init}}=N_{\mathrm{f}}$ to minimize the non-statistical bias error introduced in the SORF product approximation (14) (the “extra” input vector values were padded with zeros). This could not be done straightforwardly for SLATM, aSLATM, and (in the case of LES) SOAP since their representation vectors corresponded to $N_{\mathrm{init}}$ larger than $32768$. For SLATM, the issue was resolved by projecting all representation vectors onto linear space spawned by representation vectors of conformers in the training set, and then appending to the resulting vectors the $l^{2}$ distance between the projected and non-projected representation vector; note that using thus defined new representation vectors would not affect training and predictions of a KRR model. The procedure decreased $N_{\mathrm{init}}$ to $4096$ for SLATM, but could not be used for aSLATM and SOAP due to large number of atomic environments in the training set. Ref. [21] used Principal Component Analysis to fit atomic representations into a small enough $N_{\mathrm{init}}$, but we wanted to avoid that in order to accurately observe prediction errors of aSLATM and SOAP relative to other methods. Therefore, for aSLATM we used $N_{\mathrm{init}}=131072$, for SOAP we used $N_{\mathrm{init}}=65536$ for LES and $N_{\mathrm{init}}=32768$ for FreeSolv.
@@
We generated molecular conformers with the algorithm from Ref. [46], as implemented in the Morfeus package [47], at $T=298.15\,\mathrm{K}$ with the Merck molecular force field (MMFF94) [48, 49, 50, 51, 52, 53, 54] as implemented [55] in RDKit [56]; unimportant conformers were cut off with $\rho_{\mathrm{cut}}=0.05$. Morfeus calculations require a parameter $N_{\mathrm{conf}}^{\mathrm{Morfeus}}$ determining the number of conformers the package attempts to generate; we checked convergence of the $\mathbf{Z}\mathbf{Z}^{\mathbf{T}}$ matrix generated for one of the 4 training sets chosen for the learning curves with cMBDF representations by scanning $N_{\mathrm{conf}}^{\mathrm{Morfeus}}=8,\ldots,1024$ and observed that the values were converged at $N_{\mathrm{conf}}^{\mathrm{Morfeus}}=32$ for both LES and FreeSolv, hence $N_{\mathrm{conf}}^{\mathrm{Morfeus}}=32$ was used throughout this work. In both cases, the $\sigma$ and $\vec{\sigma}$ hyperparameters used in $\mathcal{F}$ layers were taken to be the initial guess defined in A. We used the same convergence criterion for optimizing $N_{\mathrm{tr}}$ by calculating a training set’s $\mathbf{Z}\mathbf{Z}^{\mathbf{T}}$ $N_{\mathrm{tr}}=1,2,\ldots,6,7$ for each representation and combination of MSORF layers considered. We settled on using $N_{\mathrm{tr}}=3$ throughout the work (which is also consistent with observations in Refs. [17] and [21]).
@@
Almost all representation functions used in this work were calculated using the QML2 code [57], the exception being SOAP which was calculated with DScribe package [58, 59]. While aSLATM, SLATM, FCHL19, and cMBDF were generated using default parameters present in QML2, SOAP was generated using $r_{\mathrm{cut}}^{\mathrm{SOAP}}=9.53$, $n_{\mathrm{max}}^{\mathrm{SOAP}}=8$, and $l_{\mathrm{max}}^{\mathrm{SOAP}}=8$, which were the parameters used in Ref. [60] for ML on FreeSolv with SOAP descriptors and KRR. FJK representations were generated from ab initio calculations performed with the PySCF package [61, 62, 63]. While the original paper where FJK was proposed [44] used STO-3G basis set [64, 65] and intrinsic bond orbitals [66], we found both to be ill-suited for the datasets considered in this work: Hartree-Fock STO-3G calculations sometimes failed to converge for charged molecules (which we wanted to consider as part of Slater determinant pairs), while using intrinsic bond orbitals in PySCF is complicated with some elements. Taking inspiration from previous work on using Huckel calculations for ML [7, 67], we switched to Huckel calculations with 6-311G basis [68, 69, 70, 71] and Boys localized orbitals [72] to generate FJK representations. Presence of solvent was simulated with Solvation Model based on Density (SMD) [73], an ab initio approach that has proven to be accurate in predicting experimental hydration energies [12]. When referring to “single Slater determinant” FJK we will mean representations calculated from ground state Huckel calculations for conformers in solvent (acetonitrile for LES and water for FreeSolv); for “pair Slater determinant” FJK for LES we took ground states of the molecule at charges 0 and 1 (both solvated in acetonitrile), while for FreeSolv we took ground states of the molecule in vacuum and solvated in water. For FJK we also used $l_{\mathrm{max}}=1$ and applied $\rho_{\mathrm{cut}}=0.05$ to atomic contributions to localized orbitals (see [44]).
@@
While the full procedure used to optimize hyperparameters is described in A, here we note that apart from optimizing $\lambda$ value and different $\sigma$ hyperparameters, we also considered optimizing quantity of interest by shifting it by a constant if it is intensive or by a stochiometry-dependent shift [see Eq. (29)] if it is extensive; the latter procedure was inspired by and is a more sophisticated variation of the “dressed” atom approach [74, 75]. Such optimization would mean, for example, that ML of oxidation potentials is invariant w.r.t. the choice of the reference electrode, while ML of a compound’s free energy of formation and atomization are equivalent since they only differ in reference state for different elements. While the procedure could prove useful for larger datasets, it did not improve prediction accuracy for smaller datasets used in this work, not changing the results significantly for LES and becoming numerically unstable for smaller training dataset sizes considered for FreeSolv (probably due to the large number of atom types present, leading to a large number of addtional shift-related hyperparameters). Therefore all data presented in Section 4 have been obtained without using the quantity shift procedure, while results obtained with it are left for Supplementary Data.